# Learning a Shape-adaptive Assist-as-needed Rehabilitation Policy from Therapist-informed Input

### v2

@@ IV-B Patient Motion Preferences Encoding
1: Select task and desired motion $\bm{x}_{e}$
2: Set hyperparameters NN, $L$, $C$
@@
5: Initialize parameter of linear regression function $\bm{\psi}_{T}^{0}$
@@
10:  Infer via-points $\bm{x}_{v}^{i}$ using (12)
@@
13:   Execute therapy following Section IV-A
14:   Collect $\bm{x}_{t}$ and save to $\mathcal{D}_{P}^{i}\leftarrow\mathcal{D}_{P}^{i}+\{\bm{x}_{t}\}_{t=0}^{T}$
@@
18: Output optimized parameters: $\bm{\psi}_{T}^{\star}$
The parameterized potential trajectories of patient can be modeled as
$\centering\small\bm{x}(\bm{t})=\bm{\Theta}(\bm{t})^{T}\mathbf{w}$ (6)
where $\bm{\Theta}\in\mathbb{R}^{d_{B}d_{o}\times d_{o}}$ is the basis feature matrix, $d_{B}$ is the dimensionality of the basis feature, and $\mathbf{w}\in\mathbb{R}^{d_{B}d_{o}}$ is the weight vector, assumed to follow a normal distribution $\mathbf{w}\sim\mathcal{N}(\bm{\mu}_{\mathbf{w}},\bm{\Sigma}_{\mathbf{w}})$ with unknown mean $\bm{\mu}_{\mathbf{w}}$ and covariance $\bm{\Sigma}_{\mathbf{w}}$. The parameterized probability of trajectory distribution is then represented as
where $\mathcal{P}_{p}(\bm{x}|\bm{t}_{n})$ denotes the prior probability distribution of reference trajectory estimated from the dataset $\mathcal{D}_{P}$ as shown in distribution (5). Leveraging the properties of KL-divergence for Gaussian distributions, the mean and variance can be obtained by solving two sub-problems[35]. Thanks to the kernel treatment, both mean and variance for a new input $\bm{t}^{*}$ are obtained as
$\centering\small\mathbb{E}[\bm{x}(\bm{t}^{*})]$ $=\mathbf{h}^{*}(\mathbf{H}+\lambda_{\mu}\bm{\Sigma})^{-1}\bm{\mu}$ (9)
$\mathbb{D}[\bm{x}(\bm{t}^{*})]$ $=N/\lambda_{\Sigma}(\mathbf{h}^{**}-\mathbf{h}^{*}(\mathbf{H}+\lambda_{\Sigma}\bm{\Sigma})^{-1}{\mathbf{h}^{*}}^{T})$
where $\bm{\mu}$ and $\bm{\Sigma}$ are the derived reference mean and covariance, respectively. $\mathbf{h}^{*}$, $\mathbf{h}^{**}$, and $\mathbf{H}$ denote the kernel matrices derived from the kernel function $\mathbf{h}(\cdot,\cdot;\varrho)$. $\varrho$ is the kernel parameter. $\lambda_{\mu}$ and $\lambda_{\Sigma}$ are regularization factors.
IV-C Therapist-informed Trajectory Deformation
The therapist-side robot can reproduce the actual trajectory of patient for $J$ times during each therapy iteration. The patient motion preferences can be estimated from the previous dataset $\mathcal{D}_{P}^{i}$ as in Section IV-B. For AAN therapy, the therapist applies corrective forces to partially deform the reference trajectory instead of updating the weight vector $\mathbf{w}$ to re-plan the entire reference trajectory as in (6). The therapist-applied corrective force $\bm{f}_{ext,Th}(\bm{t})$ is measured from the therapist-side robot during the reproduction of the patient’s actual trajectory. When $\bm{f}_{ext,Th}(\bm{t})$ exceeds a pretested threshold $|\overline{\bm{f}}|$, it indicates that the therapist intends to correct the therapeutic movement for the patient-side robot. As shown in Fig. 2(b), when a corrective force is activated at time $\bm{t}_{v}$, a therapist-informed via-point is inserted at $\bm{t}_{v}^{\prime}=\bm{t}_{v}+\delta\bm{t}$ to enable the partial deformation of reference trajectory. An additional dataset $\mathcal{D}_{T}$ that contains all therapist-informed via-points is introduced for one episode therapy to deform the reference trajectory for next therapy iteration. $\mathcal{D}_{T}$ shares the same structure as dataset $\mathcal{D}_{P}$ but consists of far fewer therapist-informed via-points ($L<<N$) to encourage the active participant of patients at other time.
When the therapist-informed via-points $\{\bm{t}_{l},\bm{x}_{l}^{v}\}_{l=1}^{L}$ are assumed to follow a Gaussian distribution $\mathcal{P}_{v}(\bm{x}_{l}^{v}|\bm{t}_{l})\sim\mathcal{N}(\widehat{\bm{\mu}}_{l}^{v},\widehat{\bm{\Sigma}}_{l}^{v})$, the objective in (8) for estimating $\{\bm{\mu}_{\mathbf{w}},\bm{\Sigma}_{\mathbf{w}}\}$ can be reformulated, following the idea of KMP [35], as
where $\mathcal{P}_{p}(\bm{x}|\bm{t}_{n})$ and $\mathcal{P}_{v}(\bm{x}|\bm{t}_{l})$ denote the prior probability distribution estimated from datasets $\mathcal{D}_{P}$ and $\mathcal{D}_{T}$, respectively. To achieve a closed-form solution as in (9), an extended dataset $\mathcal{D}_{U}=\mathcal{D}_{P}\cup\mathcal{D}_{T}$ is constructed, and the objective function is reformulated as
$\bm{\Sigma}(\bm{t}_{v}^{\prime})$ $=\bm{\Sigma}_{\mathbf{w}}(\bm{t}_{v}^{\prime})$
where $\beta_{\bm{\mu}}$ is the scale factor determining the amplitude of the reference trajectory deformation. Furthermore, to enable smooth complete trajectory deformation, another two via-points at $\bm{t}_{s}=0$ and $\bm{t}_{e}=\Delta T$ are inserted at the start and end points. $\Delta T$ is the duration of each therapy episode. The mean values of these via-points are defined as $\{\bm{x}_{e}(\bm{t}_{s}),\bm{x}_{e}(\bm{t}_{e})\}$ and the corresponding covariance matrices $\{\bm{\Sigma}(\bm{t}_{s}),\bm{\Sigma}(\bm{t}_{e})\}$ are defined to encourage high-precision tracking. The blue line in Fig. 2(b) represents the generated reference trajectory for the next therapy iteration, which preserves the patient’s previous motion preferences while adapting to the therapist-informed via-points.
IV-D Reproducing Therapist Input Skills
In contrast to directly re-demonstrating the reference trajectory, we constructed a latent space to represent the therapist’s input $\bm{\mathcal{\chi}}_{v}$ with the dimension of $M$. As depicted in Algorithm 1, at each ithi-th therapy iteration, the therapist state is denoted as $\bm{s}_{i}$ with a dimension of NN, defined as:
A dataset was constructed as $\mathcal{D}_{T}=\{(\bm{s}_{i},\bm{x}_{l,i}^{v})_{i=1}^{I}\}$. A partial least squares regression function is utilized to reproduce the therapist input $\bm{Y}\in\mathbb{R}^{I\times M}$ from the collected feature $\bm{X}\in\mathbb{R}^{I\times N}$, as,
$\centering\small\bm{Y}=\bm{X}\mathbf{B}$ (14)
where $\mathbf{P}$ and $\mathbf{Q}$ are the loading matrices. When only using several latent dimension, the weight matrix is derived as
where $\mathbf{W}$ is the weight matrix of the input matrix.
@@ V-A Experimental Testbed
A telerehabilitation robotic system is setup, which is taken as an example to validate the effectiveness of proposed framework for upper-limb rehabilitation. The system consists of two identical 7-DoF Franka Emika Panda robots, which are controlled by two PCs running a real-time Linux kernel at the control frequency of 1000 Hz. Both the patient-side GUI and therapist-side GUIs were developed using PyQt5 to visualize the desired and actual motions as illustrated in Fig. 1. Data transmission between the therapist-side PC and the patient-side PC was achieved using ZeroMQ within the same network, resulting in negligible communication delay. The parameters used for the lower-level interactive controllers in the subsequent rehabilitation tasks are predefined and summarized in Table I. Without loss of generality, the robot in this work is constrained to planar motion; specifically, all rehabilitation tasks are executed within the XYX–Y plane relative to the robot’s base frame (see Fig. 3).
@@
Figure 4: Performance of Task #1 by the patient at Stage #1. The first row illustrates the encoded patient motion preferences and the generated reference trajectories, while the second row shows the corrective forces applied by the therapist. (a) Iteration 1; (b) Iteration 3; (c) Iteration 6; (d) Iteration 9.
@@
V-B Experimental Protocol
The objective of the subsequent human study is to validate the effectiveness of the proposed controller in incorporating the therapist inputs for AAN therapy, instead of evaluating the biomechanical responses of the patient. Two able-bodied participants from the authors’ laboratory were recruited to play the role of therapist and patient. As illustrated in Fig. 3, participants acting as patients were unable to follow the desired motions due to the impediment introduced by the elastic band. Moreover, different levels of band elasticity were applied to simulate the patient’s various stages of recovery. Two rehabilitation tasks, each involving distinct motion patterns, are designed to demonstrate the effectiveness of the proposed controller in enabling personalized AAN therapy. For each task, one participant acts as the therapist who is assumed to be able to successfully follow the desired motion, while another participant acts as the patient to accept the AAN therapy. Therefore, given the desired motion and predefined parameters, the therapist-informed repetitive therapy will be implemented following the procedure in Algorithm 1. Two commonly used baseline methods are also implemented to demonstrate the strength of our method. First, Baseline #1 was implemented similar to [30],[5], the desired motion was selected as the reference trajectory and a variable impedance control was applied to complete the movement therapy. Second, Baseline #2 was implemented similar to [12], the therapist’s input force was directly transmitted to the patient-side robot whenever the therapist deemed assistance necessary. Furthermore, two metrics were calculated for a fair comparison. First, the corrective force measured from patient-side robot was utilized to evaluate the robot’s input and patient’s active participation, denoted as $M_{1}$. Second, movement smoothness, a commonly used indicator of kinematic control, was calculated by spectral arc length [30] to assess the quality of movement therapy, denoted as $M_{2}$.
@@ V-C Experiment Results of Human Study
The participant using the elastic band with small elasticity acts as the patient to complete two representative tasks (see Fig. 3(a)). Each task was repeated for $I=10$ therapy iterations without relying on other termination. The learned patient motion preferences and generated reference trajectories during four therapy iterations are visualized in Fig. 4 and Fig. 5.
@@ V-C1 Therapist-informed AAN Therapy
The desired motion of Task #1 (a triangle trajectory) was visualized by black lines as shown in Fig. 4. At ithi-th therapy iteration, $J=5$ therapy actual trajectories, visualized by gray lines, are collected to encode the motion preferences using GMMs with $C_{1}=10$ components. The green line and the green region represent the learned mean and variance. For fair comparison, the therapist’s skill is assumed to enable the patient to complete the key points of the desired motion. Therefore, $L_{1}=4$ via-points (see Section IV-C) were extracted from the therapist-informed corrective force and visualized by red scatters. We can see that the therapist-informed corrective force plotted in the second raw demonstrates that the therapist did not actively intervene in the current therapy session. Instead, the therapist only informed two via-points to guide the trajectory deformation for AAN therapy. The reference trajectory for next therapy iteration was then generated and shown as the dotted blue line.
@@
The desired motion of Task #2 (a rectangle trajectory) was visualized by black lines as shown in Fig. 5. At ithi-th therapy iteration, $J=5$ actual trajectories, visualized by gray lines, are collected to encode the motion preferences (green lines) using GMMs with $C_{2}=10$ components. Following the same assumption of therapist’s skills in Task #1, $L_{2}=5$ via-points were extracted from the therapist-informed corrective force visualized by red scatters. The therapist-informed corrective force plotted in the second raw indicates that the therapist only informed three via-points for patient therapy to reach three corners of the rectangle trajectory. The reference trajectory for the next therapy iteration was then generated and is illustrated by the blue line. For a fair comparison, rather than defining a specific rehabilitation termination condition, $I=10$ therapy iterations were conducted to evaluate the effectiveness of Baseline #1 and Baseline #2 on both Task #1 and Task #2 performed by the patient using the same elastic band. Baseline #1 was implemented using the variable impedance controller given the desired motion (black lines in Fig. 4 and Fig. 5) as the reference trajectory. The stiffness is modulated according to the tracking error and the damping matrix was derived from the stiffness using a fixed relation. Baseline #2 was implemented using a zero-impedance control for the patient-side robot to allow patient maximum participant, while directly adding the therapist-informed corrective force from therapist-side robot. The average values of two metrics are calculated for three therapy iterations. As shown in Fig. LABEL:Fig6:-metric1-task1 and LABEL:Fig6:-metric1-task2, Baseline #1 results in the large corrective force when the patient’s actual motion is far from the desired motion (see therapy iteration 1). Although the Baseline #2 can reduce the corrective force from the robot and encourage patient’s active engagement, as illustrated in Fig. LABEL:Fig6:-metric2-task1 and LABEL:Fig6:-metric2-task2, movement smoothness is affected by the therapist-informed therapeutic force. In summary, the proposed controller can not only maximally encourages the patient’s active participant without introducing excessive corrective force, but also ensures that the smoothness of the movement therapy remains unaffected.
@@ V-C2 Reproducing Therapist’s Therapy Skill
The collected therapist-informed via-points in dataset $\mathcal{D}_{T}$ are employed to learn regression parameters $\bm{\psi}_{T}^{\star}$ for Task #1 and Task #2. Afterwards, the via-points are generated from the regression function for patient at another recovery Stage #2, which is acted by the able-bodied participant using another elastic band with greater elasticity (see Fig. 3(b)). Results in Fig. 9(b) and Fig. 9(c) indicate that the patient can achieve the set desired motion for both Task #1 and Task #2 after $I=10$ therapy iterations without additional instruction from the therapist. The performance of reference trajectory deformation depends on the hyperparameters as shown in Table I. In this study, we assumed that the therapist primarily focused on providing corrective force to help the patient capture the key motion features enabling $L_{1}=4$ and $L_{2}=5$.
@@ VI Discussion and Conclusion
This article proposed a novel framework for therapist-in-the-loop AAN therapy. Its effectiveness and advantages in utilizing therapist’s expertise over state-of-the-art methods were validated using a telerobotics-mediated upper-limb rehabilitation system. For instance, in [18], the therapist needs to re-plan the entire reference trajectory for each therapy iteration or adapt the impedance parameters for varied therapy intensity [23]. Our framework aims to maintain patient motion preferences and partially deform the reference trajectory, which maximally encourage the active participation of patients and utilize therapist-informed corrective force. Moreover, [12],[3] rely on a low-impedance robotic system or additional sensors to detect the corrective force and add to patient-side therapy. By contrast, we build a latent space to infer the via-points for trajectory partial deformation from therapist-informed corrective force. The results demonstrate that our framework can ensure the safety by avoiding the large corrective force and mitigating the effect of communication delay for movement therapy. Most importantly, therapist-informed therapy skills can be encoded using the collected via-points in latent space and reproduced through a regression function.
This study presents a feasibility evaluation of the proposed controller with able-bodied participants under simulated motion disabilities. In future work, our framework can be extended for telerehabilitation applications by integrating the telerobotics-mediated system with advanced communication protocols (e.g., 5G). Most importantly, our framework enables that therapists only need to provide corrective force through the robot without programming the reference trajectory or manually tuning the control parameters. To further assess its therapeutic benefits, clinical studies will be conducted based on analyzing the biomechanical feedback. Currently, a partial least squares regression function for therapist’s skill reproducing is task-dependent depending on collecting labeled data from each task. Future works will focus on the in-context regression function capable of reproducing the therapeutic skill of therapist across different tasks.

### v3

@@ IV-D Reproducing Therapist Input Skills
1: Select rehabilitation task and desired motion $\bm{x}_{e}$
2: Set hyperparameters NN, $M$, $L$, $C$
@@
5: Initialize weight matrix of PLS regression function $\mathbf{B}^{0}$
@@
10:  Infer therapist-informed via-points $\bm{x}_{v}^{i}$ using (8)
@@
13:   Perform therapy following Section IV-A
14:   Collect $\bm{x}_{t}$ and store in $\mathcal{D}_{P}^{i}\leftarrow\mathcal{D}_{P}^{i}+\{\bm{x}_{t}\}_{t=0}^{T}$
@@
18: Output optimized weight matrix: $\mathbf{B}^{\star}$
@@ V-A Experimental Testbed
A telerehabilitation robotic system was set up as a testbed to validate the effectiveness of the proposed framework for upper-limb rehabilitation. The system consists of two identical 7-DoF Franka Emika Panda robots, controlled by two PCs running a real-time Linux kernel at the control frequency of 1000 Hz. Both the patient-side GUI and therapist-side GUIs were developed using PyQt5 to visualize the desired and actual motions as illustrated in Fig. 1. Data transmission between the therapist-side PC and the patient-side PC was achieved using ZeroMQ within the same network, resulting in negligible communication delay. The hyperparameters of the lower-level interactive controllers used in Algorithm 1 are predefined and summarized in Table I. Without loss of generality, the robot in this work is constrained to Cartesian space; specifically, all rehabilitation tasks are executed within the XYX–Y plane relative to the robot’s base frame (see Fig. 3).
V-B Experimental Protocol
The objective of the subsequent human study is to validate the effectiveness of the proposed controller in incorporating the therapists’ input for AAN therapy, instead of evaluating the biomechanical responses of the patient. Two able-bodied participants from the authors’ laboratory were recruited to play the roles of therapist and patient. As illustrated in Fig. 3, participants acting as patients were unable to follow the desired motions due to the impediment introduced by the elastic band. Moreover, different levels of band elasticity were applied to simulate various stages of patient recovery. Two rehabilitation tasks, each involving distinct motion patterns [8][25], were designed to demonstrate the effectiveness of the proposed controller in enabling AAN therapy. For each task, one participant acts as the therapist and is assumed to be able to accurately follow the desired motion, while another participant acts as the patient and receives AAN therapy. Accordingly, given the desired motion and predefined parameters, the therapist-informed AAN therapy was performed following the implementation in Algorithm 1. Two baseline methods were also implemented to demonstrate the strength of our method. First, Baseline #1 was implemented similar to [26],[5], the desired motion was selected as the reference trajectory and a variable impedance controller was applied to perform the movement therapy. Second, Baseline #2 was implemented similar to [11], the therapist-informed corrective force was directly transmitted to the patient-side robot whenever the therapist deems assistance necessary. Furthermore, two metrics were calculated for a fair comparison. First, the corrective force measured from patient-side robot was utilized to evaluate the robot’s input and patient’s active participation, denoted as $M_{1}$. Second, movement smoothness, a commonly used indicator of kinematic control, was calculated by spectral arc length [26] to assess the quality of movement therapy, denoted as $M_{2}$.
@@ V-B Experimental Protocol
Figure 4: Performance of Task #1 by the patient at Stage #1, where $x_{e}(t)$ denotes the desired motion, $x_{r}(t)$ represents the reference trajectory, $x_{t}(t)$ denotes the actual trajectory, $\mu_{w}(t)$ is the unknown mean of the learned motion primitive, and $\chi_{v}$ represents the therapist-informed via-point input. The first row illustrates the encoded patient motion preferences and the generated reference trajectories, while the second row shows the corrective forces applied by the therapist. (a) Iteration 1; (b) Iteration 3; (c) Iteration 6; (d) Iteration 9.
@@ V-C Experiment Results of Human Study
The participant wearing an elastic band with low elasticity acted as the patient to perform two representative tasks (see Fig. 3(a)). Each task was repeated for $I=10$ therapy iterations without relying on any termination criteria. The learned patient motion preferences and generated reference trajectories are visualized in Fig. 4 and Fig. 5.
@@ V-C1 Therapist-informed AAN Therapy
The desired motion of Task #1 (a triangular trajectory) was visualized by black lines as shown in Fig. 4. At $i$-th therapy iteration, $J=5$ actual therapy trajectories, visualized by gray lines, are collected to encode the motion preferences using GMMs with $C_{1}=10$. The green line and the shaded region represent the learned mean and variance. For a fair comparison, the therapist’s skill is assumed to enable the patient to complete the key points of the desired motion. Therefore, $L_{1}=4$ via-points (see Section IV-C) were extracted from the therapist-informed corrective force and visualized by red scatters. We can see that the therapist-informed corrective force plotted in the second row demonstrates that the therapist did not actively intervene in the current therapy session. Instead, the therapist only informed two via-points to guide the trajectory deformation for AAN therapy. The dotted blue line was the generated reference trajectory for next therapy iteration.
@@
The desired motion of Task #2 (a rectangular trajectory) was visualized by black lines as shown in Fig. 5. At $i$-th therapy iteration, $J=5$ actual therapy trajectories, visualized by gray lines, are collected to encode the motion preferences (green lines) using GMMs with $C_{2}=10$ components. Following the same assumption of therapist’s skills in Task #1, $L_{2}=5$ via-points were extracted from the therapist-informed corrective force visualized by red scatters. The therapist-informed corrective force plotted in the second row indicates that the therapist only informed three via-points for patient therapy to reach three corners of the rectangular trajectory. The reference trajectory for the next therapy iteration was then generated and is illustrated by the blue line. For a fair comparison, rather than defining a specific rehabilitation termination condition, $I=10$ therapy iterations were conducted to evaluate the effectiveness of Baseline #1 and Baseline #2 on both Task #1 and Task #2 performed by the patient using the same elastic band. Baseline #1 was implemented using the variable impedance controller given the desired motion (black lines in Fig. 4 and Fig. 5) as the reference trajectory. The stiffness is modulated according to the tracking error and the damping matrix was derived from the stiffness using a fixed relation. Baseline #2 was implemented using a zero-impedance controller for the patient-side robot to allow patient maximum participant, while directly adding the therapist-informed corrective force from therapist-side robot. The average values of two metrics are calculated for three therapy iterations. As shown in Fig. LABEL:Fig6:-metric1-task1 and LABEL:Fig6:-metric1-task2, Baseline #1 results in the large corrective force when the patient’s actual motion is far from the desired motion (see therapy iteration 1). Although the Baseline #2 can reduce the corrective force from the robot and encourage patient’s active engagement, as illustrated in Fig. LABEL:Fig6:-metric2-task1 and LABEL:Fig6:-metric2-task2, movement smoothness is affected by the therapist-informed corrective force. In summary, the proposed controller can not only maximally encourages the patient’s active participant without introducing excessive corrective force, but also ensures that the smoothness of the movement therapy remains unaffected.
@@ V-C2 Reproducing Therapist’s Therapy Skill
The collected therapist-informed via-points in dataset $\mathcal{D}_{T}$ are employed to obtain the optimal weight matrix $\mathbf{B}^{\star}$ for Task #1 and Task #2 using PLSRegression implementation from the open-source scikit-learn package. Afterwards, via-points are generated by the regression function for patient at another recovery Stage #2, which is acted by an able-bodied participant wearing another elastic band with greater elasticity (see Fig. 3(b)). Results in Fig. 9(b) and Fig. 9(c) indicate that the patient can achieve the set desired motion for both Task #1 and Task #2 after $I=10$ therapy iterations without additional instruction from the therapist. The performance of the reference trajectory deformation depends on the hyperparameters, as summarized in Table I. In this study, we assume that the therapist primarily focused on providing corrective forces to help the patient capture the key motion features with $L_{1}=4$ and $L_{2}=5$ via-points.
@@ VI Discussion and Conclusion
This article proposes a novel therapist-in-the-loop AAN therapy framework. Its effectiveness and advantages in utilizing therapist’s expertise over state-of-the-art methods were validated using a telerobotics-mediated upper-limb rehabilitation system. In [16], the therapist needs to re-plan the entire reference trajectory for each therapy iteration or to adapt impedance parameters for varied therapy intensity [7]. Our framework aims to maintain patient motion preferences while partially deforming the reference trajectory, thereby maximizing the active participation of patients and effectively incorporating therapist-informed corrective forces. Moreover, the approaches in [11],[3] rely on low-impedance robotic systems or external sensors to detect the corrective force and apply them in patient-side therapy. By contrast, we build a latent space to infer via-points for trajectory partial deformation directly from therapist-informed corrective force. The results demonstrate that our framework can ensure the safety by avoiding the large corrective force and mitigating the effect of communication delay for movement therapy. From a practical perspective, our framework enables therapists to provide corrective forces through the robot without the need to program reference trajectories or manually tune control parameters.
This study provides a feasibility evaluation of the proposed controller; however, the experimental validation is limited to two able-bodied participants under simulated motion disabilities. In future work, clinical studies will be conducted with patients at different stages of rehabilitation to further assess the therapeutic benefits of the proposed approach and to analyze the associated biomechanical responses. In this paper, the partial least squares regression model used to reproduce therapist skills is task-dependent and relies on labeled data collected for each specific task. Future work will therefore focus on developing in-context regression functions or learning methods capable of reproducing therapeutic skills across different tasks. In addition, the proposed telerobotics-mediated system will be extended to telerehabilitation applications by integrating advanced communication protocols (e.g., 5G).